the school theater sold 49 tickets to their play today. tickets for adults cost $15 and tickets for children cost $8 each. they collected a total of $469 in ticket sales today. how many of each kind of ticket did they sell?

Answer by mananth(16949)

x= number of adult tickets
y= number of child tickets
Total tickets sold
1 x + 1 y = 49 .............1

15 x + 8 y = 469 .............2
Eliminate y
multiply (1)by -8
Multiply (2) by 1
-8 x -8 y = -392
15 x + 8 y = 469
Add the two equations
7 x = 77
/ 7
x = 11
plug value of x in (1)
1 x + 1 y = 49
11 + y = 49
y = 49 -11
y = 38
y = 38
x= 11 number of adult tickets
y= 38 number of child tickets
m.ananth@hotmail.ca